[figure: 9.7 cm]

MAGNETIC PROPERTIES OF Cu(nad)_{2}(H_{2}O)_{2}SO_{4} - NEW S = 1/2

HEISENBERG ZIGZAG CHAIN?

M. Kajňaková, A. Orendáčová, M. Orendáč, J. Černák, A. Feher Faculty of Science P.J. Šafárik University, Park Angelinum 9, 041 54 Košice, Slovak

Republic

Thermodynamic properties of a polymeric compound Cu(nad)_{2}(H_{2}O)_{2}SO_{4} have been studied down to 100 mK. Since according to structural studies in the title compound the interaction between pairs of chains is mediated by hydrogen bonds, the studied material can by treated as an S = 1/2 Heisenberg zigzag spin chain. The possibility of tunning these interactions is discussed. The experimental specific heat data have been interpreted by using comparative analysis of experimental data obtained on the materials from the class Cu(X)_{2}(H_{2}O)_{2}SO_{4} (X = nad, phen, en).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Marcela Kajňaková

Address for correspondence : Faculty of Science P.J. Šafárik University, Park Angelinum 9, 041 54 Košice, Slovak Republic

Email address : mkajnak@kosice.upjs.sk